OpenStudy (anonymous):

solve using the addition method 2x+y=1 x-y=11

jimthompson5910 (jim_thompson5910):

Hint: Add the equations to get 2x+y=1 x-y=11 --------- 3x+0y =12 so, 3x = 12 solve for x, then use this to find y